# Log-majorizations between quasi-geometric type means for matrices

### v2

@@ Abstract
Keywords and phrases: $\alpha$-weighted geometric mean, $\alpha$-weighted spectral geometric mean, Rényi mean, Log-Euclidean mean, log-majorization, quantum channel, maximal $\alpha$-Rényi divergence, measured $\alpha$-Rényi divergence, $\alpha$-$z$-Rényi divergence
@@ 1 Introduction
In our previous paper [33] we considered the quasi-arithmetic matrix mean $\mathcal{A}_{\alpha,p}(A,B):=\{(1-\alpha)A^{p}+\alpha B^{p}\}^{1/p}$ and several quasi-geometric type matrix means such as the quasi $\alpha$-weighted geometric mean $G_{\alpha,p}(A,B):=(A^{p}\#_{\alpha}B^{p})^{1/p}$, the quasi versions $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ of two different $\alpha$-weighted spectral geometric means, the Rényi mean $R_{\alpha,p}$, and the Log-Euclidean mean $LE_{\alpha}$ for $0<\alpha<1$ and $p>0$ (see Section 2 for the precise definitions of these). In [33] we examined various quasi-arithmetic-geometric inequalities between each of those quasi-geometric means and $\mathcal{A}_{\alpha,q}$ with respect to different matrix orderings. In the present paper we continue to consider the above mentioned quasi-geometric type means for all $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. In fact, (1.1) is rewritten as $R_{\alpha,p}\prec_{\log}R_{\alpha,q}$, and (1.2) and (1.3) are nothing but $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$. Furthermore, the log-majorizations $G_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}G_{\alpha,p}$ were characterized in [31], and $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ was addressed in a recent paper [24]. In this paper we examine the log-majorization $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for other pairs $(\mathcal{M},\mathcal{N})$ from $R$, $G$, SGSG, $\widetilde{S}G$ and $LE$. Our goal is to hopefully obtain the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$, though we have not succeeded it for all cases. To do so, we need to address the sufficiency condition and the necessity condition separately. For the sufficiency part, since $\det\mathcal{M}_{\alpha,p}(A,B)=\det\mathcal{N}_{\alpha,q}(A,B)=(\det A)^{1-\alpha}(\det B)^{\alpha}$, the standard technique using anti-symmetric tensor powers is utilized so that the problem boils down to prove that
@@
The joint concavity/convexity of matrix trace functions in two (or more) variables has been one of major subjects in matrix analysis since Lieb’s seminal paper [42] in 1973. Lieb’s concavity [42] and Ando’s convexity [1] are especially famous, and the joint concavity/convexity of the trace functions of the forms $\mathrm{Tr}(A^{p}+B^{q})^{s}$ and $\mathrm{Tr}(A^{p/2}B^{q}A^{p/2})^{s}$ was developed in, e.g., [15, 30]. The problem of fully characterizing the joint concavity/convexity of the above latter trace function became very important, because it settles the question of the monotonicity property under CPTP maps (or the data-processing inequality) of the $\alpha$-$z$-Rényi divergence [5], as explained in detail in [14]. Then the problem was finally solved by Zhang [53] in 2020, who characterized the joint concavity/convexity of a more general trace function $(A,B)\mapsto\mathrm{Tr}(\mathrm{Tr}(A^{p/2}X^{*}B^{q}XA^{p/2})^{s}$. Also, the joint concavity of the trace function $\mathrm{Tr}(A^{p}\sigma B^{q})^{s}$ for Kubo–Ando’s operator means $\sigma$ was shown in [30]. The current status of the subject has been summarized in Carlen’s recent survey paper [12] and his new book [13]. Our second aim of this paper is to examine the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for our quasi-geometric type means $\mathcal{M}_{\alpha,p}$ and for $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$, while the case $\mathrm{Tr}\,R_{\alpha,p}$ is included in Zhang’s result in [53], and $\mathrm{Tr}\,G_{\alpha,p}$ for $0<\alpha<1$ is a special case in [30]. Our idea is to apply the relationship between the joint concavity/convexity and the monotonicity under CPTP maps for geometric type matrix functions of two variables (see Theorem 5.3) similarly to the situation in the above mentioned monotonicity question of the $\alpha$-$z$-divergence. Then as a necessary condition for joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ we have the following sandwiched inequalities (see Theorem 5.8)
@@
The structure of the paper is as follows. In Section 2 we review the definitions of the quasi-geometric matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ mentioned above as well as the notion of log-majorization. We add a few basic properties of them to those given in [33]. The main Section 3 is divided into four subsections. In those subsections we examine the log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for all pairs $(\mathcal{M},\mathcal{N})$, except for the already known cases, from $R,G,SG.\widetilde{S}G,LE$. In Section 4 we characterize the equality cases in the norm inequalities derived from the log-majorizations shown in Section 3 in terms of the commutativity of matrix variables $A,B$, expanding the former results in [28, 32]. In Section 5 we discuss the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for $\mathcal{M}=R,G,SG,\widetilde{S}G,LE$ though some cases are not new, based on theory of quantum divergences. Finally in Section 6 some concluding remarks and open problems are in order. The paper contains three appendices. Appendix A is the proof of a main part of Theorem 4.1. Appendix B is the proof of Theorem 5.3 and Appendix C is a supplement to Theorems 5.3 and 5.8.
@@ 2 Preliminaries
(2) For $0<\alpha<1$ the assertion was shown in [33, Proposition 2.2]. Let $\alpha>1$ and $A,B\geq 0$ with $s(A)\geq s(B)$. Then the proof of [33, Proposition 2.2] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,p}$ with $0<\alpha<1$ can work for any of $G_{\alpha,p}$, $R_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ with $\alpha>1$ as well. The proof for $LE_{\alpha}$ is similar to that of [36, Lemma 4.1], while we here give a short proof using [33, Lemma A.3]. Let $P_{0}:=s(B)$ and write for $0<\varepsilon<1$,
@@
Hence we can apply [33, Lemma A.3] to $Z(\varepsilon)$ with parameter $p:={1\over-\log\varepsilon}$ ($\searrow 0$ as $\varepsilon\searrow 0$) to obtain
@@
To prove the theorem we give a lemma. Note that for $0<\alpha<1$ this is contained in [33, (A.25)] where $A,B\geq 0$ are general and $P_{0}:=s(A)\wedge s(B)$.
@@
Let $L:=P_{0}(-\log A)P_{0}+P_{0}(\log B)P_{0}$ and set $Y(p):=(A^{-p/2}B^{p}A^{-p/2})^{1/p}-P_{0}e^{L}P_{0}$. Then by [33, (A.2)] we have $Y(p)\to 0$ as $p\searrow 0$. We write
@@
where the last equality follows since Taylor’s theorem (see, e.g., [29, Theorem 2.3.1]) gives
@@
with the Fréchet derivative $D(\log x)(e^{L})$ of the functional calculus by $\log x$ ($x>0$) at $e^{L}$ (see [29, p. 159]). Hence the Taylor expansion implies that $(A^{-p/2}B^{p}A^{-p/2})^{\alpha}=P_{0}+\alpha pL+o(p)$ as $p\searrow 0$. Moreover, note that $A^{p}\#_{\alpha}B^{p}=P_{0}(A^{p}\#_{\alpha}B^{p})P_{0}$. This is obvious for $0<\alpha<1$ (even for general $A,B\geq 0$). For $\alpha>1$ this can be seen from
@@
For $0<\alpha<1$ the assertion was shown in [33, Theorem 2.3]. For $\alpha>1$ let $A,B\geq 0$ with $s(A)\geq s(B)$. For $G_{\alpha,p}$ Lemma 2.3 gives $G_{\alpha,p}(A,B)^{p}=P_{0}+pK+o(p)$, where $K:=(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. This implies that with $\mathcal{H}_{0}$ being the range of $P_{0}$,
@@
showing the result for $G_{\alpha,p}$. For $R_{\alpha,p}$ apply (2.9) to $A^{\alpha-1}$ and $B^{\alpha}$ in place of $A,B$ to have $A^{{1-\alpha\over 2}p}B^{\alpha p}A^{{1-\alpha\over 2}p}=P_{0}+pK+o(p)$. This implies that $\log R_{\alpha,p}(A,B)\big|_{\mathcal{H}_{0}}=(K+o(1))|_{\mathcal{H}_{0}}$, showing the result for $R_{\alpha,p}$. The proof for $SG_{\alpha,p}$ is similar to that of [33, Theorem 2.3] for $0<\alpha<1$. Finally, for $\widetilde{S}G_{\alpha,p}$ by Lemma 2.3 we have $A^{-p}\#_{\alpha}B^{p}=P_{0}+p\widetilde{K}+o(p)$, where $\widetilde{K}:=(1-\alpha)P_{0}(-\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. Therefore,
@@
and the log-majorization $X\prec_{\log}Y$ means that $X\prec_{w\log}Y$ and $\prod_{i=1}^{n}\lambda_{i}(X)=\prod_{i=1}^{n}\lambda_{i}(Y)$, i.e., $\det X=\det Y$. Details on (weak) log-majorization are found in [2], [10, Chap. II] and [43]. Some basic properties of several matrix orderings including $\leq,\leq_{\lambda},\prec_{(w)\log}$ mentioned above are also found in [33, Sec. 2.2].
@@
(1) Let $(\mathcal{M},\mathcal{N})$ be any pair from $G,SG,\widetilde{S}G,R,LE$, and let $\alpha,\beta\in(0,\infty)\setminus\{1\}$ and $p,q>0$. When $\alpha\neq\beta$, $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\beta,q}$ are not definitively comparable even for positive scalar variables, as explained in [33, Remark 2.7(1)]. So we only compare between $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\alpha,q}$ under the same weight parameter $\alpha$.
@@
(4) For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ except for the three cases mentioned in (3), note that there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\mathcal{M}_{\alpha,p}(A,B)\not\leq_{\lambda}\mathcal{N}_{\alpha,q}(A,B)$. Indeed, assume that this is not the case. Then by (2) we must have $\lambda(\mathcal{M}_{\alpha,p}(A,B))=\lambda(\mathcal{N}_{\alpha,q}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$. But this fails to hold except for the above three cases and the trivial case of $\mathcal{M}=\mathcal{N}$ and $p=q$.
@@
We can easily compute all the expressions in the lemma by applying [33, Lemma 3.6] to the relevant situations discussed in [33]. See [33, Lemmas 3.4, 3.5 and 4.17] (with $y=x$) for (2.15), (2.11) and (2.12) respectively, and see the proofs of [33, Theorems 4.29 and 4.37] for (2.13) and (2.14) respectively, while details of computations are left to the reader. Also we note that all the relevant expressions in [33] for $0<\alpha<1$ remain valid for all $\alpha\in(0,\infty)\setminus\{1\}$. ∎
To verify Remark 2.4(4) more explicitly, we here include brief discussions for completeness. Consider $A_{0},B_{\theta}\in\mathbb{M}_{2}^{++}$ as in Lemma 2.5. First let $(\mathcal{M}_{\alpha,p},\mathcal{M}_{\alpha,q})$ be be given for $\mathcal{M}\in\{R,G,SG,\widetilde{S}G\}$. Assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{M}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$, which implies from (2.11)–(2.15) that ${-1-x^{p}+x^{\alpha p}+x^{(1-\alpha)p}\over p(1-x^{p})}={-1-x^{q}+x^{\alpha q}+x^{(1-\alpha)q}\over q(1-x^{q})}$ etc. Letting $x\searrow 0$ gives $p=q$ in each case. Next let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be given for $\mathcal{M}\neq\mathcal{N}$, and assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{N}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$. When $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})=(R_{\alpha,p},G_{\alpha,q})$, by (2.11) and (2.12) we have

### v3

@@ Abstract
Keywords: Log-majorization, Quasi matrix means, Weighted geometric mean, Weighted spectral geometric mean, Rényi mean, Log-Euclidean mean, $\alpha$-Rényi divergences
@@ 1 Introduction
In our previous paper [35] we considered the quasi-arithmetic matrix mean $\mathcal{A}_{\alpha,p}(A,B):=\{(1-\alpha)A^{p}+\alpha B^{p}\}^{1/p}$ and several quasi-geometric type matrix means such as the quasi $\alpha$-weighted geometric mean $G_{\alpha,p}(A,B):=(A^{p}\#_{\alpha}B^{p})^{1/p}$, the quasi versions $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ of two different $\alpha$-weighted spectral geometric means, the Rényi mean $R_{\alpha,p}$, and the Log-Euclidean mean $LE_{\alpha}$ for $0<\alpha<1$ and $p>0$ (see Section 2 for the precise definitions of these). In [35] we examined various quasi-arithmetic-geometric inequalities between each of those quasi-geometric means and $\mathcal{A}_{\alpha,q}$ with respect to different matrix orderings. In the present paper we continue to consider the above mentioned quasi-geometric type means for all $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$. In fact, (1.1) is rewritten as $R_{\alpha,p}\prec_{\log}R_{\alpha,q}$, and (1.2) and (1.3) are nothing but $G_{\alpha,p}\prec_{\log}G_{\alpha,q}$. Furthermore, the log-majorizations $G_{\alpha,p}\prec_{\log}R_{\alpha,q}$ and $R_{\alpha,q}\prec_{\log}G_{\alpha,p}$ were characterized in [33], and $R_{\alpha,q}(A,B)\prec_{\log}SG_{\alpha,p}(A,B)$ was addressed in a recent paper [26]. In this paper we examine the log-majorization $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for other pairs $(\mathcal{M},\mathcal{N})$ from $\{R,G,SG,\widetilde{S}G,LE\}$. Our goal is to hopefully obtain the necessary and sufficient condition on $p,q,\alpha$ under which $\mathcal{M}_{\alpha,p}(A,B)\prec_{\log}\mathcal{N}_{\alpha,q}(A,B)$ holds for all $A,B>0$, though we have not succeeded it for all cases. To do so, we need to address the sufficiency condition and the necessity condition separately. For the sufficiency part, since $\det\mathcal{M}_{\alpha,p}(A,B)=\det\mathcal{N}_{\alpha,q}(A,B)=(\det A)^{1-\alpha}(\det B)^{\alpha}$, the standard technique using anti-symmetric tensor powers is utilized so that the problem boils down to prove that
@@
The joint concavity/convexity of matrix trace functions in two (or more) variables has been one of major subjects in matrix analysis since Lieb’s seminal paper [45] in 1973. Lieb’s concavity [45] and Ando’s convexity [1] are especially famous, and the joint concavity/convexity of the trace functions of the forms $\mathrm{Tr}(A^{p}+B^{q})^{s}$ and $\mathrm{Tr}(A^{p/2}B^{q}A^{p/2})^{s}$ was developed in, e.g., [15, 32]. The problem of fully characterizing the joint concavity/convexity of the above latter trace function became very important, because it settles the question of the monotonicity property under CPTP maps (or the data-processing inequality) of the $\alpha$-$z$-Rényi divergence [5], as explained in detail in [14]. Then the problem was finally solved by Zhang [57] in 2020, who characterized the joint concavity/convexity of a more general trace function $(A,B)\mapsto\mathrm{Tr}(A^{p/2}X^{*}B^{q}XA^{p/2})^{s}$. Also, the joint concavity of the trace function $\mathrm{Tr}(A^{p}\sigma B^{q})^{s}$ for Kubo–Ando’s operator means $\sigma$ [43] was shown in [32]. The current status of the subject has been summarized in Carlen’s recent survey paper [12] and his new book [13]. Our second aim of this paper is to examine the joint concavity/convexity of the trace functions $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for our quasi-geometric type means $\mathcal{M}_{\alpha,p}$ and for $\alpha\in(0,\infty)\setminus\{1\}$ and $p>0$, while the case $\mathrm{Tr}\,R_{\alpha,p}$ is included in Zhang’s result in [57], and $\mathrm{Tr}\,G_{\alpha,p}$ for $0<\alpha<1$ is a special case in [32]. Our idea is to apply the relationship between the joint concavity/convexity and the monotonicity under CPTP maps for geometric type matrix functions of two variables (see Theorem 5.3) similarly to the situation in the above mentioned monotonicity question of the $\alpha$-$z$-divergence. Then as a necessary condition for joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ we have the following sandwiched inequalities (see Theorem 5.8)
@@
The structure of the paper is as follows. In Section 2 we review the definitions of the quasi-geometric matrix means $R_{\alpha,p}$, $G_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ and $LE_{\alpha}$ mentioned above as well as the notion of log-majorization. We add a few basic properties of them to those given in [35]. The main Section 3 is divided into four subsections. In those subsections we examine the log-majorizations $\mathcal{M}_{\alpha,p}\prec_{\log}\mathcal{N}_{\alpha,q}$ for all pairs $(\mathcal{M},\mathcal{N})$, except for the already known cases, from $\{R,G,SG,\widetilde{S}G,LE\}$. In Section 4 we characterize the equality cases in the norm inequalities derived from the log-majorizations shown in Section 3 in terms of the commutativity of matrix variables $A,B$, expanding the former results in [30, 34]. In Section 5 we discuss the joint concavity/convexity of $\mathrm{Tr}\,\mathcal{M}_{\alpha,p}$ for $\mathcal{M}=R,G,SG,\widetilde{S}G,LE$ though some cases are not new, based on theory of quantum divergences. Finally in Section 6 some concluding remarks and open problems are in order. The paper contains three appendices. Appendix A is the proof of a main part of Theorem 4.1. Appendix B is the proof of Theorem 5.3 and Appendix C is a supplement to Theorems 5.3 and 5.8.
@@ 2 Preliminaries
(2) For $0<\alpha<1$ the assertion was shown in [35, Proposition 2.2]. Let $\alpha>1$ and $A,B\geq 0$ with $s(A)\geq s(B)$. Then the proof of [35, Proposition 2.2] for $SG_{\alpha,p}$ and $\widetilde{S}G_{\alpha,p}$ with $0<\alpha<1$ can work for any of $G_{\alpha,p}$, $R_{\alpha,p}$, $SG_{\alpha,p}$, $\widetilde{S}G_{\alpha,p}$ with $\alpha>1$ as well. The proof for $LE_{\alpha}$ is similar to that of [38, Lemma 4.1], while we here give a short proof using [35, Lemma A.3]. Let $P_{0}:=s(B)$ and write for $0<\varepsilon<1$,
@@
Hence we can apply [35, Lemma A.3] to $Z(\varepsilon)$ with parameter $p:={1\over-\log\varepsilon}$ ($\searrow 0$ as $\varepsilon\searrow 0$) to obtain
@@
To prove the theorem we give a lemma. Note that for $0<\alpha<1$ this is contained in [35, (A.25)] where $A,B\geq 0$ are general and $P_{0}:=s(A)\wedge s(B)$.
@@
Let $L:=P_{0}(-\log A)P_{0}+P_{0}(\log B)P_{0}$ and set $Y(p):=(A^{-p/2}B^{p}A^{-p/2})^{1/p}-P_{0}e^{L}P_{0}$. Then by [35, (A.2)] we have $Y(p)\to 0$ as $p\searrow 0$. We write
@@
where the last equality follows since Taylor’s theorem (see, e.g., [31, Theorem 2.3.1]) gives
@@
with the Fréchet derivative $D(\log x)(e^{L})$ of the functional calculus by $\log x$ ($x>0$) at $e^{L}$ (see [31, p. 159]). Hence the Taylor expansion implies that $(A^{-p/2}B^{p}A^{-p/2})^{\alpha}=P_{0}+\alpha pL+o(p)$ as $p\searrow 0$. Moreover, note that $A^{p}\#_{\alpha}B^{p}=P_{0}(A^{p}\#_{\alpha}B^{p})P_{0}$. This is obvious for $0<\alpha<1$ (even for general $A,B\geq 0$). For $\alpha>1$ this can be seen from
@@
For $0<\alpha<1$ the assertion was shown in [35, Theorem 2.3]. For $\alpha>1$ let $A,B\geq 0$ with $s(A)\geq s(B)$. For $G_{\alpha,p}$ Lemma 2.3 gives $G_{\alpha,p}(A,B)^{p}=P_{0}+pK+o(p)$, where $K:=(1-\alpha)P_{0}(\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. This implies that with $\mathcal{H}_{0}$ being the range of $P_{0}$,
@@
showing the result for $G_{\alpha,p}$. For $R_{\alpha,p}$ apply (2.9) to $A^{\alpha-1}$ and $B^{\alpha}$ in place of $A,B$ to have $A^{{1-\alpha\over 2}p}B^{\alpha p}A^{{1-\alpha\over 2}p}=P_{0}+pK+o(p)$. This implies that $\log R_{\alpha,p}(A,B)\big|_{\mathcal{H}_{0}}=(K+o(1))|_{\mathcal{H}_{0}}$, showing the result for $R_{\alpha,p}$. The proof for $SG_{\alpha,p}$ is similar to that of [35, Theorem 2.3] for $0<\alpha<1$. Finally, for $\widetilde{S}G_{\alpha,p}$ by Lemma 2.3 we have $A^{-p}\#_{\alpha}B^{p}=P_{0}+p\widetilde{K}+o(p)$, where $\widetilde{K}:=(1-\alpha)P_{0}(-\log A)P_{0}+\alpha P_{0}(\log B)P_{0}$. Therefore,
@@
and the log-majorization $X\prec_{\log}Y$ means that $X\prec_{w\log}Y$ and $\prod_{i=1}^{n}\lambda_{i}(X)=\prod_{i=1}^{n}\lambda_{i}(Y)$, i.e., $\det X=\det Y$. Details on (weak) log-majorization are found in [2], [10, Chap. II] and [46]. Some basic properties of several matrix orderings including $\leq,\leq_{\lambda},\prec_{(w)\log}$ mentioned above are also found in [35, Sec. 2.2].
@@
(1) Let $(\mathcal{M},\mathcal{N})$ be any pair from $\{R,G,SG,\widetilde{S}G,LE\}$, and let $\alpha,\beta\in(0,\infty)\setminus\{1\}$ and $p,q>0$. When $\alpha\neq\beta$, $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\beta,q}$ are not definitively comparable even for positive scalar variables, as explained in [35, Remark 2.7(1)]. So we only compare between $\mathcal{M}_{\alpha,p}$ and $\mathcal{N}_{\alpha,q}$ under the same weight parameter $\alpha$.
@@
[…]
(5) For any pair $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ except for the three cases mentioned in (3), note that there exist $A,B\in\mathbb{M}_{2}^{++}$ such that $\mathcal{M}_{\alpha,p}(A,B)\not\leq_{\lambda}\mathcal{N}_{\alpha,q}(A,B)$. Indeed, assume that this is not the case. Then by (2) we must have $\lambda(\mathcal{M}_{\alpha,p}(A,B))=\lambda(\mathcal{N}_{\alpha,q}(A,B))$ for all $A,B\in\mathbb{M}_{2}^{++}$. But, as verified below, this fails to hold except for the above three cases and the trivial case of $\mathcal{M}=\mathcal{N}$ and $p=q$.
@@
We can easily compute all the expressions in the lemma by applying [35, Lemma 3.6] to the relevant situations discussed in [35]. See [35, Lemmas 3.4, 3.5 and 4.17] (with $y=x$) for (2.15), (2.11) and (2.12) respectively, and see the proofs of [35, Theorems 4.29 and 4.37] for (2.13) and (2.14) respectively, while details of computations are left to the reader. Also we note that all the relevant expressions in [35] for $0<\alpha<1$ remain valid for all $\alpha\in(0,\infty)\setminus\{1\}$. ∎
To verify Remark 2.4(5) more explicitly, we here include brief discussions for completeness. Consider $A_{0},B_{\theta}\in\mathbb{M}_{2}^{++}$ as in Lemma 2.5. First let $(\mathcal{M}_{\alpha,p},\mathcal{M}_{\alpha,q})$ be given for $\mathcal{M}\in\{R,G,SG,\widetilde{S}G\}$. Assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{M}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$, which implies from (2.11)–(2.15) that ${-1-x^{p}+x^{\alpha p}+x^{(1-\alpha)p}\over p(1-x^{p})}={-1-x^{q}+x^{\alpha q}+x^{(1-\alpha)q}\over q(1-x^{q})}$ etc. Letting $x\searrow 0$ gives $p=q$ in each case. Next let $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})$ be given for $\mathcal{M}\neq\mathcal{N}$, and assume that $\lambda_{1}(\mathcal{M}_{\alpha,p}(A_{0},B_{\theta}))=\lambda_{1}(\mathcal{N}_{\alpha,q}(A_{0},B_{\theta}))$ for all $x\in(0,1)$. When $(\mathcal{M}_{\alpha,p},\mathcal{N}_{\alpha,q})=(R_{\alpha,p},G_{\alpha,q})$, by (2.11) and (2.12) we have